# Chameleon: Automated Color Palette Adaptation for Dark Mode Data Visualizations

### v1

@@ 3.5. Simulated-Annealing-based Optimization
Generating a new color palette (Lines 7–9). At each iteration, the algorithm generates a new palette $P_{\text{new}}$ by randomly adjusting one color in $P_{\text{current}}$. We operate in the LCH (Lightness, Chroma, Hue) color space, which provides more intuitive control over color properties than RGB (schanda2007colorimetry). For each adjustment, we randomly select one of three color components with equal probability: lightness (L), chroma (C), or hue (H). If lightness is selected, we adjust its value by a random amount between -20 and +20 units and ensure the result stays within [0, 100], where 0 represents black and 100 represents white. For chroma adjustments, we similarly apply a random change between -20 and +20 units, constrained to [0, 100], where higher values indicate more color saturation. If hue is selected, we modify it by a random value between -50 and +50 degrees and constrain it to [0, 360], representing the full color wheel. Since not all colors in LCH color space can be represented on digital displays that use sRGB color space, we verify that each perturbed color falls within the valid sRGB gamut by checking whether its transformed RGB values lie between 0 and 1. If a perturbed color is outside this gamut, the algorithm continues generating new perturbations until a valid, displayable color is found or a maximum attempt limit is reached. This controlled random perturbation strategy helps the algorithm explore the color space effectively while ensuring all generated colors can be correctly displayed.
@@ 4.1.1. Comparisons on Color Palettes
Chameleon’s automated transformations closely resemble Apple’s predefined dark mode palette. Apple’s iOS design guidelines (appleColorApple) offers a small set of predefined pairs of colors for light and dark modes, where each light mode color corresponds to a specific dark mode counterpart (e.g., light red becomes dark red). Using Apple’s predefined light mode colors (first row in Figure 4) as input, we applied Chameleon to transform these colors into dark mode equivalents. Each light mode color (e.g., [Uncaptioned image]) was processed through our algorithm, generating its corresponding dark mode version (e.g., [Uncaptioned image]). We tested eight different colors, and the transformed dark mode colors are displayed in Figure 4. From the top-to-bottom comparison in the last row of Figure 4, we see that the dark mode colors generated by Chameleon closely match the predefined dark mode colors from Apple (the second row of Figure 4). However, the iOS predefined color palette is limited in the number of color pairs available, making it less adaptable for visualizations with more color requirements. In contrast, Chameleon’s automated transformation process allows it to scale to a larger variety of colors, enabling use across visualizations that may exceed the limited set provided by iOS.
@@ 4.1.2. Visualization Examples from Dark Mode Guideline Documentation
Our method transforms and produces results that align with the official dark mode design guidelines, such as those provided by Android. As seen in Figure 2[Uncaptioned image] and Figure 2[Uncaptioned image], the Android dark mode design document (androidDarkMode) includes a pair of visualization examples for light and dark modes. Comparing Figure 2[Uncaptioned image] (official dark mode example) and Figure 2[Uncaptioned image] (output from Chameleon), we observe similarities. Notably, the colors of the bar chart in our dark mode transformation closely match those provided in the official Android design. Additionally, text elements were automatically converted from black to white to maintain legibility against the dark background.
@@ 4.2. Expert Interview
To understand how professionals would use a tool like Chameleon in their workflow, and how they would perceive its performance, we conducted semi-structured interviews with four visualization design and UI design professionals. As shown in Table 1, two of the experts (E1 and E2) have over four years of experience in visualization design research. E1 specializes in visual analysis system design, while E2 focuses on color design in visualizations. The other two experts (E3 and E4) have over one year of industry experience, with E3 working as a visualization designer for a newspaper and E4 as a UI/UX designer. Each interview lasted approximately 60 minutes and began with a brief introduction of the project. We then presented several cases for discussion, allowing participants to share their insights and comments. The interviews were recorded and transcribed using automated speech-to-text technology (Radford2022RobustSR), followed by a thematic analysis to summarize their feedback.
@@ 4.3.1. Evaluation Design
We evaluated 85 charts from the D3 Graph Gallery (d3graphgalleryGraphGallery), spanning 37 distinct visualization types from basic (bar, line, pie charts) to complex forms (streamgraphs, choropleth maps). This diverse sample allowed us to test our method across varied visualization types and color palettes. We compared three conditions:
@@ 4.3.2. Metrics and Procedures
We assessed two key metrics: Legibility measured the contrast ratio between visual elements and background according to WCAG 2.1 AA guidelines (w3ContentAccessibility), requiring a minimum 3:1 ratio:
@@ 4.4.1. Study Design
We evaluated three visualization conditions: light mode, dark mode, and inverse mode. The study comprised three types of tasks. For legibility assessment, participants rated how clearly they could read single charts on a 7-point Likert scale (1=very difficult, 7=very easy). In the color similarity assessment, participants compared how similar dark mode and inverse mode visualizations were to their light mode versions (7-point scale, 1=very different, 7=very similar). For analytical tasks, participants performed identification and comparison tasks on multiple-view visualizations, following Munzner’s task categorization (munzner2014visualization). We measured accuracy, completion time, and visual fatigue (via a 10-question survey (Heuer1989RestPO)).
@@
For the analytical tasks, participants performed value identification and comparison tasks on multiple-view visualizations. Each visualization was paired with two questions: one focused on identifying values and the other on comparing them. This framework was based on Tamara Munzner’s task categorization (munzner2014visualization), which classifies analytical tasks into identification, comparison, and summary. The identification and comparison tasks (Table 2) were selected because they are fundamental analytic tasks and require users to engage intensively with the visualizations. Tasks were designed to evaluate user performance across the three visualization modes. Metrics included task accuracy, completion time, and a visual fatigue index derived from a 10-question survey (Heuer1989RestPO), with higher scores indicating more fatigue.
@@ 4.4.2. Materials
We used two sets of visualizations. The first set included 16 single charts from the D3 Graph Gallery, covering a range of common visualization types (e.g., line charts, bar charts, pie charts) and complex types (e.g., choropleth maps, streamgraphs). The second set comprised 4 multiple-view visualizations from Tableau Public Gallery (tableauGallery), divided into two categories (Figure 9): visualizations where colors had conventional meanings (e.g., green for increase, red for decrease in stock price) and visualizations where colors had no inherent meaning.
@@ 5. Discussion
Strengths and Limitations. The evaluation of Chameleon demonstrates its effectiveness in transforming light mode visualizations while maintaining color semantics and legibility. However, several limitations emerged: Chameleon treats all visualization elements equally, not distinguishing between text, axes, and graphical components that require different contrast levels according to WCAG 2.1 guidelines (w3ContentAccessibility). This limitation, identified in expert feedback, often led to poor contrast and legibility in text and axes. In future, a post-processing step to detect and invert text and axis colors could mitigate this issue. Some color choices, particularly greens and yellows, were considered visually unappealing due to prioritizing legibility over aesthetics. Additionally, Chameleon assumes similar color interactions in both modes and depends on the input visualization’s contrast quality. The tool’s relatively long execution time, while not critical for offline processing, could limit real-time applications.
@@ 6. Conclusion and Future Work
In future work, we plan to broaden the usability of Chameleon by extending support beyond static images to include scalable vector graphics (SVG), enabling finer control over individual elements. We also plan to adapt the method for mobile app interfaces, where dynamic context (e.g., changing light conditions) requires adaptive and lightweight transformations. Together, these directions will enhance the practicality of Chameleon for real-world applications and improve accessibility across devices and platforms. Further, with the increasing usage of Large Language Models (LLMs) for visualization tasks (fan2025well; wang2023llm4vis), it will also be interesting to explore how LLMs can benefit automated color palette adaptation for dark mode visualizations.
@@ Acknowledgements.
This project is supported by the Ministry of Education, Singapore, under its Academic Research Fund Tier 2 (Proposal ID: T2EP20222-004).

### v2

@@ 3.5. Simulated-Annealing-based Optimization
Generating a new color palette (Lines 7–9). At each iteration, the algorithm generates a new palette $P_{\text{new}}$ by randomly adjusting one color in $P_{\text{current}}$. We operate in the LCH (Lightness, Chroma, Hue) color space, which provides more intuitive control over color properties than RGB (Schanda, 2007). For each adjustment, we randomly select one of three color components with equal probability: lightness (L), chroma (C), or hue (H). If lightness is selected, we adjust its value by a random amount between -20 and +20 units and ensure the result stays within [0, 100], where 0 represents black and 100 represents white. For chroma adjustments, we similarly apply a random change between -20 and +20 units, constrained to [0, 100], where higher values indicate more color saturation. If hue is selected, we modify it by a random value between -50 and +50 degrees and constrain it to [0, 360], representing the full color wheel. Since not all colors in LCH color space can be represented on digital displays that use sRGB color space, we verify that each perturbed color falls within the valid sRGB gamut by checking whether its transformed RGB values lie between 0 and 1. If a perturbed color is outside this gamut, the algorithm continues generating new perturbations until a valid, displayable color is found or a maximum attempt limit is reached. This controlled random perturbation strategy helps the algorithm explore the color space effectively while ensuring all generated colors can be correctly displayed.
@@ 4.1.1. Comparisons on Color Palettes
Chameleon’s automated transformations closely resemble Apple’s predefined dark mode palette. Apple’s iOS design guidelines (Apple, 2025a) offers a small set of predefined pairs of colors for light and dark modes, where each light mode color corresponds to a specific dark mode counterpart (e.g., light red becomes dark red). Using Apple’s predefined light mode colors (first row in Figure 4) as input, we applied Chameleon to transform these colors into dark mode equivalents. Each light mode color (e.g., [Uncaptioned image]) was processed through our algorithm, generating its corresponding dark mode version (e.g., [Uncaptioned image]). We tested eight different colors, and the transformed dark mode colors are displayed in Figure 4. From the top-to-bottom comparison in the last row of Figure 4, we see that the dark mode colors generated by Chameleon closely match the predefined dark mode colors from Apple (the second row of Figure 4). However, the iOS predefined color palette is limited in the number of color pairs available, making it less adaptable for visualizations with more color requirements. In contrast, Chameleon’s automated transformation process allows it to scale to a larger variety of colors, enabling use across visualizations that may exceed the limited set provided by iOS.
@@ 4.1.2. Visualization Examples from Dark Mode Guideline Documentation
Our method transforms and produces results that align with the official dark mode design guidelines, such as those provided by Android. As seen in Figure 2[Uncaptioned image] and Figure 2[Uncaptioned image], the Android dark mode design document (Material Design (2025), Google) includes a pair of visualization examples for light and dark modes. Comparing Figure 2[Uncaptioned image] (official dark mode example) and Figure 2[Uncaptioned image] (output from Chameleon), we observe similarities. Notably, the colors of the bar chart in our dark mode transformation closely match those provided in the official Android design. Additionally, text elements were automatically converted from black to white to maintain legibility against the dark background.
@@ 4.2. Expert Interview
To understand how professionals would use a tool like Chameleon in their workflow, and how they would perceive its performance, we conducted semi-structured interviews with four visualization design and UI design professionals. As shown in Table 1, two of the experts (E1 and E2) have over four years of experience in visualization design research. E1 specializes in visual analysis system design, while E2 focuses on color design in visualizations. The other two experts (E3 and E4) have over one year of industry experience, with E3 working as a visualization designer for a newspaper and E4 as a UI/UX designer. Each interview lasted approximately 60 minutes and began with a brief introduction of the project. We then presented several cases for discussion, allowing participants to share their insights and comments. The interviews were recorded and transcribed using automated speech-to-text technology (Radford et al., 2023), followed by a thematic analysis to summarize their feedback.
@@ 4.3.1. Evaluation Design
We evaluated 85 charts from the D3 Graph Gallery (Holtz, 2025), spanning 37 distinct visualization types from basic (bar, line, pie charts) to complex forms (streamgraphs, choropleth maps). This diverse sample allowed us to test our method across varied visualization types and color palettes. We compared three conditions:
@@ 4.3.2. Metrics and Procedures
We assessed two key metrics: Legibility measured the contrast ratio between visual elements and background according to WCAG 2.1 AA guidelines (World Wide Web Consortium, 2025), requiring a minimum 3:1 ratio:
@@ 4.4.1. Study Design
We evaluated three visualization conditions: light mode, dark mode, and inverse mode. The study comprised three types of tasks. For legibility assessment, participants rated how clearly they could read single charts on a 7-point Likert scale (1=very difficult, 7=very easy). In the color similarity assessment, participants compared how similar dark mode and inverse mode visualizations were to their light mode versions (7-point scale, 1=very different, 7=very similar). For analytical tasks, participants performed identification and comparison tasks on multiple-view visualizations, following Munzner’s task categorization (Munzner, 2025). We measured accuracy, completion time, and visual fatigue (via a 10-question survey (Heuer et al., 1989)).
@@
For the analytical tasks, participants performed value identification and comparison tasks on multiple-view visualizations. Each visualization was paired with two questions: one focused on identifying values and the other on comparing them. This framework was based on Tamara Munzner’s task categorization (Munzner, 2025), which classifies analytical tasks into identification, comparison, and summary. The identification and comparison tasks (Table 2) were selected because they are fundamental analytic tasks and require users to engage intensively with the visualizations. Tasks were designed to evaluate user performance across the three visualization modes. Metrics included task accuracy, completion time, and a visual fatigue index derived from a 10-question survey (Heuer et al., 1989), with higher scores indicating more fatigue.
@@ 4.4.2. Materials
We used two sets of visualizations. The first set included 16 single charts from the D3 Graph Gallery, covering a range of common visualization types (e.g., line charts, bar charts, pie charts) and complex types (e.g., choropleth maps, streamgraphs). The second set comprised 4 multiple-view visualizations from Tableau Public Gallery (Tableau, 2025), divided into two categories (Figure 9): visualizations where colors had conventional meanings (e.g., green for increase, red for decrease in stock price) and visualizations where colors had no inherent meaning.
@@ 5. Discussion
Strengths and Limitations. The evaluation of Chameleon demonstrates its effectiveness in transforming light mode visualizations while maintaining color semantics and legibility. However, several limitations emerged: Chameleon treats all visualization elements equally, not distinguishing between text, axes, and graphical components that require different contrast levels according to WCAG 2.1 guidelines (World Wide Web Consortium, 2025). This limitation, identified in expert feedback, often led to poor contrast and legibility in text and axes. In future, a post-processing step to detect and invert text and axis colors could mitigate this issue. Some color choices, particularly greens and yellows, were considered visually unappealing due to prioritizing legibility over aesthetics. Additionally, Chameleon assumes similar color interactions in both modes and depends on the input visualization’s contrast quality. The tool’s relatively long execution time, while not critical for offline processing, could limit real-time applications.
@@ 6. Conclusion and Future Work
In future work, we plan to broaden the usability of Chameleon by extending support beyond static images to include scalable vector graphics (SVG), enabling finer control over individual elements. We also plan to adapt the method for mobile app interfaces, where dynamic context (e.g., changing light conditions) requires adaptive and lightweight transformations. Together, these directions will enhance the practicality of Chameleon for real-world applications and improve accessibility across devices and platforms. Further, with the increasing usage of Large Language Models (LLMs) for visualization tasks (Fan et al., 2025; Wang et al., 2023), it will also be interesting to explore how LLMs can benefit automated color palette adaptation for dark mode visualizations.
@@ Acknowledgements.
This project is supported by the Ministry of Education, Singapore, under its Academic Research Fund Tier 2 (Proposal ID: T2EP20222-0049).